SW patch message daemons and masterdiagnostic

Staff:

LBTO: Xianyu,Guido,Juan Carlos,Gustavo
Arcetri: Fabio and Alfio.

Summary.

Tersting the new version of the WFS/Adsec and AOS software. Software deployed and tested without problems. Close loop test done for all the Bin modes. It has been reported some issues for the pupil centering at Bin3 and bin4.
Software roll back to default and wait for Saturday E/D night to be deployed and tested on sky.

Details.

Telescope release for AO SW at.8:30am

Setting up the telescope.

Software version updated for AOS, Adsec and WFS.
Power ON WFS.
Gave problems during the power on of the OCAM.

Tried several times Power ON and the OCAM come alive. Investigation.

Alfio detected that WFS was deployed in a different brach. Software compiled.

Roll back software/telescope configuration for the internet lost to the mountain.

  • 9:am Mountain lost internet connection.

9:45 Reconfigure for daytime tests.

9:45 internet back. Green light from the Telescope to continue with the AO SW tests.

Call the mountain and we can start to prepare sw/telescope for the tests.
Fabio called and preparing the SW.
9:55. Argos Cal Unit sent to working position and it is not moving. Calling Gustavo.

Reset and it didn't work. suggestion to move the telescope to 45. Asking mountain crew to tilt the telescope to 55 degrees.
Mountain crew in a meeting and we asked to tilt the telescope from our side.Not problems to tilt and back to zenith.

At 85 didn't triggered the park:YES. Moving to 45 and the flag was triggered at 55.

BAck to zenith and stow pins in and deploy the Argos cal Unit.


settings for various binnings:

Bin1 - ARGOS shutter = full open freq 1.7kHz

Bin2 - ARGOS shutter =63, R13.2 -> freq 1000Hz (TO BE UPDATED AFTER BULB replacement)

Bin3 - ARGOS shutter = 50, WFS mag 15.6

Bin4 - ARGOS shutter = 45, WFS mag 17.1 -> freq 280Hz

10:47 going thru preset, failed at optimize gain

10:49 adding some disturbance to help the gain optimize, seeing 1”

10:50 preset AO failed…

10:51 opened loop manually and preset AO again

10:54 optimize gain

10:56 applying optical gain (soul.source)

10:57 taking Optical loop gain 20210312_175952

11:00 taking Optical loop gain 20210312_180020

11:01 taking Optical loop gain 20210312_180058

11:01 cancel preset

11:02 loop open

11:02 (bin2) shutter 62 to have magnitude 13

11:04 changing disturbance to work at 1kHz

11:10 optimize gain, plots are displayed only for HO2 and not for the other mode group (GUI stuck)

11:13 taking Optical loop gain 20210312_181252

11:13 (bin3) shutter 49 to have magnitude 15.6

11:14 changing disturbance to work at 500Hz

11:15 pupil centering seems to not converge

11:19 loop opened…

11:25 preset

11:27 all lights off in the dome

11:27 centering pupil

11:27 pupil centering not converging

11:30 timeout caused the AO loop to stop

11:30 cancel preset

11:35 preset AO

11:36 acquire ref. (intervention mode)

11:42 center pupils

11:43 check flux

11:43 close loop (camera frequency 829 Hz)

11:47 close loop did not converge, we forced to open loop

11:47 (bin4) shutter 45 to have magnitude 17.1

11:49 preset AO

11:50 FLAO GUI is not responding…

12:06 close loop (camera frequency 437 Hz)

12:09 we are not sure what is the loop status

12:09 taking Optical loop gain 20210312_191142

12:09 taking Optical loop gain 20210312_191425

12:15 open loop

12:15 WFS power off

12:15 turn on dome lights

12:17 rest shell

12:18 power off

12:22 restarted software with old adsec version

12:24 set and rest....

NOTE:

The Chamber was not complete dark, so for the faint AO mode could have an impact in the system behaviour (spurious lights...)

emails report:

Xianyu email report:

The AO SW test (the SOUL-devel with MasterDiagnostic and MessageDaemon
on both sxadsec and soul-sxwfs) was successful this morning. Please see
the detailed information below:

For bin1 we successfully went through all the steps, including gain
optimization and optical gain application.
For bin2, we successfully went through all the steps, including gain
optimization and optical gain application.
For bin3, we successfully closed the loop, but the camera lens cannot
catch up during gain optimization and not try optical gain application.
For bin4, we successfully closed the loop, including gain optimization
but failed with optical gain application.
Since we were running out of time and not software issue, we did not fix
the camera lens issues for bin3 and bin4.

The test was running with 1” disturbance seeing, please the detailed
loop information and track numbers below:

Bin1 - ARGOS shutter = full open, freq 1.7kHz, TN:20210312_175952,
20210312_180020 and 20210312_180058
Bin2 - ARGOS shutter =62, R13.2, freq 1KHz, TN:20210312_181252
Bin3 - ARGOS shutter =49, R15.6, freq 500Hz, TN: none
Bin4 - ARGOS shutter = 45, R17.1, freq 437Hz, TN:20210312_191142,
20210312_191425

Thanks to Gustavo for his help with ARGOS Swing arm issues.

Please let us if you have anymore questions.

Cheers,
Alfio,Fabio,Guido, Juan Carlos and Xianyu

P.S., after the test, both sxadsec and soul-sxwfs are revert original
version. During the test, we also installed the new AOS version with
more DIMM info, and kept it running after the test.

Fabio email report:

Ciao Barry and Greg,

today with Xianyu and Guido we tested the sw version which includes fixes for Message Daemon and Master Diagnostic bugs (both on adsec and wfs). The test went well and the next step would be to test it in nighttime too, but off course it is up to you to decide when to do it. Currently the old version is running.

I take it from Xianyu you already have the instructions to use the old or the new version on the adsec. Regarding the wfs things are similar, with some addendum. At the end of this email you can find the instructions, however tomorrow I am available around this time to leave the system in the state you plan to use it in the night.

Ciao

Fabio

From old to new version:

log into soul-sxwfs

w_stop_all

cd /home/AOeng/SOUL.latest

source flao_environment.sh

sudo su

cp /etc/init/keepUp.conf.new /etc/init/keepUp.conf

exit

wfs_init

# all the GUIs must be then launched from this terminal or another where you did:

cd /home/AOeng/SOUL.latest

source flao_environment.sh

From new to old version:

log into soul-sxwfs

cd /home/AOeng/SOUL.latest

source flao_environment.sh

w_stop_all

sudo su

cp /etc/init/keepUp.conf.old /etc/init/keepUp.conf

exit

cd /home/AOeng/SOUL

source flao_environment.sh

wfs_init

-- JuanCarlosGuerra - 12 Mar 2021
Topic revision: r3 - 12 Mar 2021, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback